### v3

@@ 3.2.5 Phase 5: Data Synthesis
During this phase, we synthesized the extracted data that we gathered from the previous phase (i.e., Phase 4: Data extraction) in order to answer the five RQs (see Table 1). We used open coding & constant comparison to analyze several data items and answer certain RQs of this SMS (see Table 4). Open coding & constant comparison are two widely used techniques in software engineering (Seaman, 1999) during qualitative data analysis. Open coding generates codes for incidents that can be further classified into concepts and categories (Seaman, 1999). Constant comparison is a continuous process for verifying the generated concepts and categories. Both concepts and categories evolve and saturate until they fit the data (Seaman, 1999). Specifically, in this SMS, open coding & constant comparison were employed for analyzing the data items D5 and D6 to answer RQ2 (sources), the data items D9 and D9 to answer RQ4 (approaches and tools), and the data item D10 to answer RQ5 (challenges). In addition, descriptive statistics (Wohlin et al., 2003) was also used for analyzing the data items D5 and D6 to answer RQ2 (sources), the data items D8 and D9 to answer RQ4 (approaches and tools), and the data item D10 to answer RQ5 (challenges). Descriptive statistics can provide quantitative summaries based on the initial description of the extracted data. Moreover, we used a hybrid approach which is a combination of open coding & constant comparison, predefined classifications (i.e., the conceptual model for architectural description in the ISO 42010:2011 standard (ISO/IEC/IEEE, 2011), the categories of architecture decisions by Kruchten (2004), and the categories of architectural changes by Williams and Carver (2010)), and descriptive statistics for analyzing the data item D4 to answer RQ1 (mined architectural information). On another hand, a hybrid approach which is a combination of open coding & constant comparison, predefined classifications of architecting activities by Hofmeister et al. (2007), Tang et al. (2010), and Li et al. (2013), and descriptive statistics was employed for analyzing the data item D7 to answer RQ3 (architecting activities).
@@
As mentioned above, we utilized a qualitative data analysis tool MAXQDA to support the analysis process. MAXQDA assists human annotators to label text segments within their contexts and assign them to categories. Specifically, before the formal data analysis (manual labeling), to reach an agreement about the data items that we gathered from the previous phase (i.e., Phase 4: Data extraction), we first performed a pilot data analysis. Specifically, this analysis process involved the following steps: (1) The first and second authors independently checked and read a random sample of 5 primary studies (from 104 selected studies, see Figure 2). (2) The first two authors independently labeled the extracted data with codes that succinctly summarize the data items (see Table 4) for answering RQs. (3) The first two authors independently grouped all the codes into higher-level concepts and turned them into categories or subcategories. The grouping process was iterative, in which each author continuously went back and forth between the code, concepts, categories, and extracted data items to revise and refine both the code, concepts, and categories. In order to improve the reliability of the pilot data analysis results, the first two authors held a meeting and followed a negotiated agreement approach (Campbell et al., 2013) to compare the coding results, then discussed their disagreements and uncertain judgments on the data encoding results in an effort to reconcile them and arrive at a final version of the pilot data analysis results in which all the discrepancies have been resolved. The first author carried on with the formal data analysis and followed the same steps used during the pilot data analysis. In the following paragraph, we provide the details of the formal data analysis process.
@@
The first author fully read the full text of each selected study from the remaining primary studies. Subsequently, he summarized the main ideas stated in several sentences in the selected study. He continued to encode the summarized ideas (from each selected study) to generate codes, concepts, and categories and subcategories. For example, when answering RQ1 (mined architectural information), he referred to these sentences in the selected study [S27] that states: “(…) Implementation of a tactic often spreads across more than one source file. Consequently, domain topics that motivate an architectural tactic are often not fully presented in the tactical file itself but also appear in neighboring files (i.e., files which use tactical files or are used by tactical files). Therefore, we need to identify the whole context in which a tactic is implemented. The context analysis is comprised of two steps. First, we identify complete tactic instances whose implementation spreads across multiple tactical files. Second, we consider the full technical context of an implemented tactic instance (…) To discover this full context in which a tactic instance is implemented, we use Algorithm 2 to discover the full context in which each tactic instance is implemented (…)”. In this case, he summarized and encoded those sentences as the code “identify the whole context in which a tactic is implemented”. Afterwards, the first author grouped this code into a higher-level concept (i.e., “tactic and context”). Then, he applied constant comparison to compare the concepts identified in one summarized idea with the concepts that emerged from other summarized ideas to identify the concepts which have similar semantic meanings. He proceeded to group similar concepts into main categories and subcategories. For example, the concept “tactic and context” was merged into the subcategory “architectural tactic and context relationship” that was further categorized into the category “design relationship” (see Table 5). It is worth noting that some primary studies (e.g., [S47]) may not explicitly clarify the concrete categories of mined architectural information. For such cases, if a study claims that it mines architectural information without providing a concrete type of this information, we thoroughly checked the content of that study including examples or excerpts from software repositories (if any) that show the mined architectural artifacts, as well as, the dataset of the study in the replication package (if available). Doing so, we tried to examine and categorize the data for identifying the concrete types of mined architectural information. Moreover, as stated before in this section, we used predefined categorization of architectural information to generate categories and subcategories of mined architecture information for answering RQ1. Among others, we utilized an existing category of architectural design decisions by Kruchten (2004) (i.e., structural, behavioral, and ban decisions) to classify the mined architectural decisions. To mitigate the personal bias during the formal data analysis, the second and third authors of this SMS reviewed and validated the generated codes, concepts, categories, and subcategories. The results of this SMS are provided in Section 4.
@@ 4 Results
We first present the overview of the selected studies in Section 4.1. The results of five RQs are reported in Section 4.2 to Section 4.6.
@@ 4.1 Overview of the Selected Studies
The selected studies were published in 55 venues. We only list the venues that publish more than one selected studies (see Table 10 in Appendix B). Note that, merged conferences were only counted as one conference, for example, Working International Conference on Software Architecture (WICSA) and International Conference on Software Architecture (ICSA) were counted as one conference. As shown in Table 10, some venues tend to publish more studies on mining architectural information. For instance, the dedicated conferences on software architecture record the most publications (i.e., 17), which are International Conference on Software Architecture (ICSA) and European Conference on Software Architecture (ECSA) with 10 and 7 publications, respectively. The conference with the most number of studies (i.e., 10 studies) is ICSA. In terms of journals, both the Journal of Systems and Software and IEEE Transactions on Software Engineering index the most studies with 6 and 5 studies, respectively. As shown in Figure 4, the 104 studies cover three types of publication venues, i.e., conference, journal, and workshop. Most of the selected studies were published in conferences (i.e., 63.5%, 66 out of 104) compared with journals (i.e., 28.8%, 30 out of 104) and workshops (i.e., 7.7%, 8 out of 104), which indicates that conferences are the popular venues to disseminate the work on mining architectural information.
@@ 4.2 Results of RQ1: Mined Architectural Information
As described in Section 3.2.5, a hybrid approach (which is a combination of open coding & constant comparison, predefined classifications, and descriptive statistics) was used to analyze the extracted data item D4 (see Table 4) and identify the mined architectural information from software repositories. Our data synthesis generated 7 main categories and 29 subcategories of mined architectural information (see Table 5), in which architectural description (69.2%, 72 out of 104 studies), architectural solution (42.3%, 44 out of 104 studies), architectural decision (30.8%, 32 out of 104 studies), system requirement (29.8%, 31 out of 104 studies) are the top three most mined architectural information to support the development process. Note that, one study (e.g., [S1]) may utilize an approach to mine more than one type of architectural information and this affects the accuracy of the statistical representation (i.e., the total percentage is more than 100% in Table 5).
Category Subcategory Studies Count
Architectural description (69.2%, 72 out of 104)
Architectural model
[S11] [S12] [S29] [S30] [S36] [S37] [S39] [S46] [S49] [S51] [S54] [S62] [S64] [S67] [S68] [S73] [S75] [S80] [S83] [S86]
20
Architectural view
[S8] [S16] [S18] [S30] [S35] [S37] [S39] [S46] [S47] [S54] [S59] [S61] [S64] [S68] [S98] [S70] [S81] [S102]
18
Architectural rationale
[S1] [S2] [S4] [S11] [S18] [S23] [S24] [S25] [S26] [S28] [S38] [S47] [S79]
13
System of interest
[S11] [S26] [S29] [S39] [S60] [S72] [S85] [S100] [S103]
9
Architectural concern
[S4] [S6] [S7] [S8] [S9] [S11] [S38] [S66]
8
Architectural solution (42.3%, 44 out of 104)
Architectural pattern
[S1] [S2] [S3] [S4] [S16] [S23] [S24] [S28] [S32] [S44] [S49] [S50] [S52] [S53] [S54] [S61] [S63] [S69] [S70] [S82] [S79]
21
Architectural tactic
[S1] [S4] [S6] [S7] [S13] [S26] [S9] [S24] [S49] [S50] [S53] [S56] [S57] [S79]
14
Framework
[S3] [S4] [S14] [S16] [S26] [S49] [S50] [S52] [S61]
9
Architectural decision (30.8%, 32 out of 104)
Technology decision
[S2] [S3] [S4] [S14] [S15] [S16] [S18] [S23] [S33] [S38] [S53] [S85]
12
Structural decision
[S17] [S18] [S19] [S22] [S23] [S32] [S33] [S47] [S54] [S85]
10
Behavioral decision
[S18] [S19] [S22] [S33] [S42] [S47] [S54]
7
Ban or non-existence decision
[S19] [S22] [S33]
3
System requirement (29.8%, 31 out of 104)
Quality attribute
[S1] [S2] [S3] [S4] [S8] [S11] [S16] [S17] [S23] [S24] [S26] [S28] [S32] [S42] [S61] [S66] [S76] [S83]
18
Functional requirement
[S1] [S4] [S17] [S43] [S45] [S54] [S55] [S83] [S79]
9
Constraint
[S1] [S2] [S3] [S54]
4
Architectural technical debt (23.0%, 24 out of 104)
Architectural smell
[S5] [S29] [S65] [S71] [S84] [S88] [S90] [S92] [S93] [S94] [S97] [S98] [S99] [S100]
14
Architectural anti-pattern
[S37] [S60] [S64] [S72] [S95] [S104]
6
Architectural compliance issue
[S30] [S34] [S36] [S73]
4
Architectural change (17.3%, 18 out of 104)
Perfective change
[S17] [S18] [S32] [S35] [S40] [S11] [S77]
7
Corrective change
[S35] [S32] [S40] [S46] [S77]
5
Adaptive change
[S35] [S40] [S77]
3
Preventative change
[S35] [S40] [S77]
3
Design relationship (16.3%, 17 out of 104)
Component-Component relationship
[S12] [S20] [S48] [S59] [S67] [S69] [S74] [S101] [S103]
9
Architectural tactic-Context relationship
[S6] [S27]
2
Architectural pattern-Architectural pattern relationship
[S63] [S78]
2
Architectural pattern-Context relationship
[S87]
1
Architectural pattern-Quality attribute relationship
[S31]
1
Architectural tactic-Quality attribute relationship
[S6]
1
Design pattern-Architectural tactic relationship
[S21]
1
Architectural description is one of the important sources used for communicating and sharing architectural information about the high-level design of a software system (Clements et al., 2003). For example, the information in architectural description, such as architectural views, which describe systems in multiple views (e.g., logical view, development view), enable the architecture to be communicated and understood by various stakeholders (e.g., architects, developers, and project managers) (Kruchten, 1995). 69.2% of the selected studies (72 out of 104) proposed approaches and tools for mining or recovering architectural descriptions of software systems from several artifacts (e.g., code [S29][S39][S46][S51]) that are contained in various software repositories, such as VCS (e.g., GitHub), to assist the development process. We further classified the mined architectural information in the architectural description category by using the conceptual model for architectural description in the ISO 42010:2011 standard (ISO/IEC/IEEE, 2011), which defines a set of architectural elements that make up the architectural description model (ISO/IEC/IEEE, 2011). As shown in Table 5, we collected five subcategories of mined architectural information (e.g., architectural model, architectural view) in the architectural description category.
Architectural model is a partial abstraction of a system. Architectural model is normally a diagram that can be created by using specific notations, such as Unified Model Language (UML), in which the primary aim is to illustrate a specific set of architectural design elements (e.g., components) inherent in the structure and design of a system. The architecture modeling process can be bottom-up, by which details of the system are built utilizing knowledge about components and interconnections and how they compose together to realize the characteristics of the system. Alternatively, it can be top-down, by which details of the components and interconnections are extracted from knowledge of the whole system (Mistrík et al., 2014). Architectural model is the most mined (20 out of 72 studies) architectural information in architectural description category. For example, Granchelli et al. [S39] developed a semi-automatic approach and tool (i.e., MicroART) based on Model-Driven Engineering (MDE) and Domain-Specific Language (DSL) for automatically recovering architectural models of microservice-based systems from source code. The proposed approach can also allow an architect to manually refine the initial extracted architecture model into refined architectural models that are suitable for his/her needs, such as for performing architecture change impact analysis. Figure 5(a) shows an example of mined architectural model from a microservice-based system collected from GitHub.
Architectural view represents the system from the viewpoint of different stakeholders, such as developers, system analysts, project managers, and end-users. The architect chooses and develops a set of views that will enable the architecture to be communicated to, and understood by, the stakeholders of the system. In capturing or representing the system in architectural views, the architect will typically create one or more architecture models (in the form of diagrams), possibly using specific notations (e.g., UML). An architecture view comprises selected parts of the system, chosen so as to demonstrate to a particular stakeholder or group of stakeholders that their concerns are being adequately addressed in the architecture. 18 out of 104 studies proposed approaches and tools for mining architectural views to support various architecture design tasks. For example, in [S46] an architecture recovery approach and a prototype tool (named ARCADE) were developed to help software engineers (e.g., architects) automatically extract and recover various architectural views (including development views) from source code. The mined architectural views can assist in determining, for example, when, how, and to what extent system-level or component-level architectural views evolve during software development. Figure 5(b) depicts an example of mined architectural view (i.e., development view) from a software system.
Architectural rationale records explanation, justification, and reasoning about architectural design decisions that are made in software architecture design. Lopez et al. [S24] presented an ontology-based approach named TREx (Toeska Rationale Extraction) for extracting, representing, and exploring architectural rationale information from text documents (such as email, wiki, and meeting notes). Specifically, this approach consists of three components: (1) pattern-based information extraction to recover rationale, (2) ontology-based representation of rationale and architectural concepts, and (3) facet-based interactive exploration of rationale. Initial results from applying TREx suggest that architectural rationale (such as reasons behind architectural decisions) can be semi-automatically extracted and mined from a project’s unstructured text documents. The mined architectural rationale provides several benefits in software development, such as rationale reuse, rationale auditing, and architect training. Figure 5(c) shows an example of architectural rationale mined from a Q&A site (i.e., Stack Overflow).
System of interest denotes a specific type of a system (such as a microservice-based system) whose architecture is under exploration during the development. Some studies proposed approaches and tools for mining architectural information from specific types of systems. For example, Soldani et al. in [S29] proposed an approach and toolchain (i.e., $\mu$TOSCA) for mining and refactoring architectural smells in microservice-based systems.
@@
(a) Architectural model mined from source code of a microservice-based system ([S39])
(b) Architectural view (i.e., development view) mined from a software system ([S81])
(c) Architectural rationale mined from the Stack Overflow site ([S4])
@@
Architectural solutions are the fundamental building blocks in architecture design, and they are used to address architecture design concerns (Bass et al., 2012). 44 out of 104 studies (42.3%) researched on mining architectural solutions to support the development. We identified three subcategories of architectural solutions that are mined from repositories, namely architectural pattern, architectural tactic, and framework (see Table 5). For example, Liu et al. [S63] utilized a semi-automatic approach for mining architectural pattern use scenarios and related pattern pairs from specific software development blogs and Q&A site (i.e., Stack Overflow). This mined architectural information can aid developers to acquaint with the usage and relatedness of architecture patterns in today’s modern software development. On the other hand, Chinnappan et al. [S9] employed a manual approach (i.e., qualitative data analysis) to mine energy-aware architectural tactics (such as energy savings mode tactic, stop current task & recharge tactic) for Robotic Operating System (ROS) based software systems from GitHub, ROS Answers, and Stack Overflow. To do so, they carried out a multi-phase study that resulted in seven energy-awareness architectural tactics. To foster the applicability of the identified tactics even beyond the ROS community, they described these tactics in a generic implementation-independent manner by utilizing diagrams inspired by UML component and sequence diagram notations. These energy-aware architectural tactics can serve as guidance for roboticists, as well as other developers interested in architecting and implementing energy-aware software. Then again, Márquez et al. [S26] proposed a semi-automatic approach for mining frameworks in OSS for addressing scalability concerns in microservice applications, and these frameworks can assist developers in designing scalable microservice applications.
@@
Architectural decision: Mining and understanding architectural decisions can help inform future architectural decisions and implementation choices, and can avoid introducing architectural inefficiencies (Shahbazian et al., 2018) or architectural decay later (Hassaine et al., 2012). Thus, 32 out of 104 studies (30.8%) focused on mining architectural decisions to assist the development. We categorized the mined architectural decisions into four subcategories (including structural decision, behavioral decision, and technology decision) by following existing classifications of architectural design decisions in (Kruchten, 2004) and (Bass et al., 2003) (see Table 5). For example, Bhat et al. [S22] employed an ML-based approach to mine and classify architectural decisions from issue tracking systems (e.g., Jira) into three subcategories, namely structural, behavioral, and non-existence decisions. The mined architectural decisions in this category can help architects and developers utilize these decisions in similar development contexts. Figure 6(a) illustrates an example of mined architectural decisions (i.e., structure decisions) from an issue tracking system.
@@
System requirement: Analyzing requirements to relate and map them to their corresponding architectural elements (e.g., architectural components) has become one of the major challenges faced by architects during the development process (Casamayor et al., 2012a). The failure of a high percentage of software projects is often caused by, for example, the lack of proper requirements analysis, incomplete requirement specifications, and changing requirements, among others (Hull et al., 2005). 31 out of 104 studies (29.8%) focus on mining the requirements of systems from repositories to assist architecting process (see Table 5). We further categorized the mined system requirements into three subcategories: Quality Attributes (QAs), functional requirement, and constraint by following the classification described in (Cervantes and Kazman, 2016). For instance, in [S66], Gokyer et al. used an approach based on NLP and ML techniques for automatically mining Non-Functional Requirements (NFRs) expressed in plain text and mapping these NFRs to architectural elements, such as architectural components, in the problem domain. The mined architectural information can guide architects in making architectural decisions effectively. On the other hand, in [S43], Casamayor et al. presented an approach based on NLP and ML techniques for semi-automatically mining and analyzing functional requirements (from the textual description of requirements) that will become the responsibilities of certain architectural components in the system, in order to help bridge the gap between requirements analysis and architectural design. Then again Soliman et al. [S3] developed a Web-based search engine for searching several types of architectural information (including constraints) from Q&A sites.

### v4

@@ 3.2.5 Phase 5: Data Synthesis
During this phase, we synthesized the extracted data that we gathered from the previous phase (i.e., Phase 4: Data extraction) in order to answer the five RQs (see Table 1). We used open coding & constant comparison to analyze several data items and answer certain RQs of this SMS (see Table 4). Specifically, in this SMS, open coding & constant comparison were employed for analyzing the data items D5 and D6 to answer RQ2 (sources), the data items D9 and D9 to answer RQ4 (approaches and tools), and the data item D10 to answer RQ5 (challenges). In addition, descriptive statistics (Wohlin et al., 2003) was also used for analyzing the data items D5 and D6 to answer RQ2 (sources), the data items D8 and D9 to answer RQ4 (approaches and tools), and the data item D10 to answer RQ5 (challenges). Descriptive statistics can provide quantitative summaries based on the initial description of the extracted data. Moreover, we used a hybrid approach which is a combination of open coding & constant comparison, predefined classifications (i.e., the conceptual model for architectural description in the ISO 42010:2011 standard (ISO/IEC/IEEE, 2011), the categories of architecture decisions by Kruchten (2004), and the categories of architectural changes by Williams and Carver (2010)), and descriptive statistics for analyzing the data item D4 to answer RQ1 (mined architectural information). On another hand, a hybrid approach which is a combination of open coding & constant comparison, predefined classifications of architecting activities by Hofmeister et al. (2007), Tang et al. (2010), and Li et al. (2013), and descriptive statistics was employed for analyzing the data item D7 to answer RQ3 (architecting activities).
@@
As mentioned above, we utilized a qualitative data analysis tool MAXQDA [^1] https://www.maxqda.com/ to support the analysis process. Specifically, before the formal data analysis (manual labeling), to reach an agreement about the data items that we gathered from the previous phase (i.e., Phase 4: Data extraction), we first performed a pilot data analysis. Specifically, this analysis process involved the following steps: (1) The first and second authors independently checked and read a random sample of 5 primary studies (from 104 selected studies, see Figure 2). (2) The first two authors independently labeled the extracted data with codes that succinctly summarize the data items (see Table 4) for answering RQs. (3) The first two authors independently grouped all the codes into higher-level concepts and turned them into categories or subcategories. The grouping process was iterative, in which each author continuously went back and forth between the code, concepts, categories, and extracted data items to revise and refine both the code, concepts, and categories. In order to improve the reliability of the pilot data analysis results, the first two authors held a meeting and followed a negotiated agreement approach (Campbell et al., 2013) to compare the coding results, then discussed their disagreements and uncertain judgments on the data encoding results in an effort to reconcile them and arrive at a final version of the pilot data analysis results in which all the discrepancies have been resolved. The first author carried on with the formal data analysis and followed the same steps used during the pilot data analysis. In the following paragraph, we provide the details of the formal data analysis process.
@@
The first author fully read the full text of each selected study from the remaining primary studies. Subsequently, he summarized the main ideas stated in several sentences in the selected study. He continued to encode the summarized ideas (from each selected study) to generate codes, concepts, and categories and subcategories. For example, when answering RQ1 (mined architectural information), he referred to these sentences in the selected study [S27] that state: “(…) Implementation of a tactic often spreads across more than one source file. Consequently, domain topics that motivate an architectural tactic are often not fully presented in the tactical file itself but also appear in neighboring files (i.e., files which use tactical files or are used by tactical files). Therefore, we need to identify the whole context in which a tactic is implemented (…) In this case, he summarized and encoded those sentences as the code “identify the whole context in which a tactic is implemented”. Afterwards, the first author grouped this code into a higher-level concept (i.e., “tactic and context”). Then, he applied constant comparison to compare the concepts identified in one summarized idea with the concepts that emerged from other summarized ideas to identify the concepts which have similar semantic meanings. He proceeded to group similar concepts into main categories and subcategories. For example, the concept “tactic and context” was merged into the subcategory “architectural tactic and context relationship” that was further categorized into the category “design relationship” (see Table LABEL:minedArchitecturalInfo). It is worth noting that some primary studies (e.g., [S47]) may not explicitly clarify the concrete categories of mined architectural information. For such cases, if a study claims that it mines architectural information without providing a concrete type of this information, we thoroughly checked the content of that study including examples or excerpts from software repositories (if any) that show the mined architectural artifacts, as well as, the dataset of the study in the replication package (if available). Doing so, we tried to examine and categorize the data for identifying the concrete types of mined architectural information. To mitigate the personal bias during the formal data analysis, the second and third authors of this SMS reviewed and validated the generated codes, concepts, categories, and subcategories. The results of this SMS are provided in Section 4.
@@ 4.1 Overview of the Selected Studies
The selected studies were published in 55 venues. We only list the venues that publish more than one selected studies (see the Publication Venues in the Supplementary Material (de Dieu et al., 2024)). Note that, merged conferences were only counted as one conference, for example, Working International Conference on Software Architecture (WICSA) and International Conference on Software Architecture (ICSA) were counted as one conference. As shown in the Publication Venues in de Dieu et al. (2024), some venues tend to publish more studies on mining architectural information. For instance, the dedicated conferences on software architecture record the most publications (i.e., 17), which are International Conference on Software Architecture (ICSA) and European Conference on Software Architecture (ECSA) with 10 and 7 publications, respectively. The conference with the most number of studies (i.e., 10 studies) is ICSA. In terms of journals, both the Journal of Systems and Software and IEEE Transactions on Software Engineering index the most studies with 6 and 5 studies, respectively. As shown in Figure 4, the 104 studies cover three types of publication venues, i.e., conference, journal, and workshop. Most of the selected studies were published in conferences (i.e., 63.5%, 66 out of 104) compared with journals (i.e., 28.8%, 30 out of 104) and workshops (i.e., 7.7%, 8 out of 104), which indicates that conferences are the popular venues to disseminate the work on mining architectural information.
@@ 4.2 Results of RQ1: Mined Architectural Information
As described in Section 3.2.5, a hybrid approach (which is a combination of open coding & constant comparison, predefined classifications, and descriptive statistics) was used to analyze the extracted data item D4 (see Table 4) and identify the mined architectural information from software repositories. Our data synthesis generated 7 main categories and 29 subcategories of mined architectural information (see Table LABEL:minedArchitecturalInfo), in which architectural description (69.2%, 72 out of 104 studies), architectural solution (42.3%, 44 out of 104 studies), architectural decision (30.8%, 32 out of 104 studies), system requirement (29.8%, 31 out of 104 studies) are the top three most mined architectural information to support the development process. Note that, one study (e.g., [S1]) may utilize an approach to mine more than one type of architectural information and this affects the accuracy of the statistical representation (i.e., the total percentage is more than 100% in Table LABEL:minedArchitecturalInfo).
Architectural description is one of the important sources used for communicating and sharing architectural information about the high-level design of a software system (Clements et al., 2003). For example, the information in architectural description, such as architectural views, which describe systems in multiple views (e.g., logical view, development view), enable the architecture to be communicated and understood by various stakeholders (e.g., architects, developers, and project managers) (Kruchten, 1995). 69.2% of the selected studies (72 out of 104) proposed approaches and tools for mining or recovering architectural descriptions of software systems from several artifacts (e.g., code [S29][S39][S46][S51]) that are contained in various software repositories, such as VCS (e.g., GitHub), to assist the development process. We further classified the mined architectural information in the architectural description category by using the conceptual model for architectural description in the ISO 42010:2011 standard (ISO/IEC/IEEE, 2011), which defines a set of architectural elements that make up the architectural description model (ISO/IEC/IEEE, 2011). As shown in Table LABEL:minedArchitecturalInfo, we collected five subcategories of mined architectural information (e.g., architectural model, architectural view) in the architectural description category.
Architectural model is a partial abstraction of a system. Architectural model is normally a diagram that can be created by using specific notations, such as Unified Model Language (UML), in which the primary aim is to illustrate a specific set of architectural design elements (e.g., components) inherent in the structure and design of a system. The architecture modeling process can be bottom-up, by which details of the system are built utilizing knowledge about components and interconnections and how they compose together to realize the characteristics of the system. Alternatively, it can be top-down, by which details of the components and interconnections are extracted from knowledge of the whole system. Architectural model is the most mined (20 out of 72 studies) architectural information in architectural description category. For example, Granchelli et al. [S39] developed a semi-automatic approach and tool (i.e., MicroART) based on Model-Driven Engineering (MDE) and Domain-Specific Language (DSL) for automatically recovering architectural models of microservice-based systems from source code. The proposed approach can also allow an architect to manually refine the initial extracted architecture model into refined architectural models that are suitable for his/her needs, such as for performing architecture change impact analysis.
Architectural view represents the system from the viewpoint of different stakeholders, such as developers, system analysts, project managers, and end-users. The architect chooses and develops a set of views that will enable the architecture to be communicated to, and understood by, the stakeholders of the system. In capturing or representing the system in architectural views, the architect will typically create one or more architecture models (in the form of diagrams), possibly using specific notations (e.g., UML). An architecture view comprises selected parts of the system, chosen to demonstrate to a particular stakeholder or group of stakeholders that their concerns are being adequately addressed in the architecture. 18 out of 104 studies proposed approaches and tools for mining architectural views to support various architecture design tasks. For example, in [S46] an architecture recovery approach and a prototype tool (named ARCADE) were developed to help software engineers (e.g., architects) automatically extract and recover various architectural views (including development views) from source code. The mined architectural views can assist in determining, for example, when, how, and to what extent system-level or component-level architectural views evolve during software development. Figure 5(a) depicts an example of mined architectural view (i.e., development view) from a software system.
Architectural rationale records explanation, justification, and reasoning about architectural design decisions that are made in software architecture design. Lopez et al. [S24] presented an ontology-based approach named TREx (Toeska Rationale Extraction) for extracting, representing, and exploring architectural rationale information from text documents (such as email, wiki, and meeting notes). Specifically, this approach consists of three components: (1) pattern-based information extraction to recover rationale, (2) ontology-based representation of rationale and architectural concepts, and (3) facet-based interactive exploration of rationale. Initial results from applying TREx suggest that architectural rationale (such as reasons behind architectural decisions) can be semi-automatically extracted and mined from a project’s unstructured text documents. The mined architectural rationale provides several benefits in software development, such as rationale reuse, rationale auditing, and architect training. Figure 5(b) shows an example of architectural rationale mined from a Q&A site (i.e., Stack Overflow).
System of interest denotes a specific type of a system (such as a microservice system) whose architecture is under exploration during the development. Some studies proposed approaches and tools for mining architectural information from specific types of systems. For example, Soldani et al. in [S29] proposed an approach and toolchain (i.e., $\mu$TOSCA) for mining and refactoring architectural smells in microservice systems.
@@
(a) Architectural view (i.e., development view) mined from a software system ([S81])
(b) Architectural rationale mined from the Stack Overflow site ([S4])
@@
Architectural solutions are the fundamental building blocks in architecture design, and they are used to address architecture design concerns (Bass et al., 2012). 44 out of 104 studies (42.3%) researched on mining architectural solutions to support the development. We identified three subcategories of architectural solutions that are mined from repositories, namely architectural pattern, architectural tactic, and framework (see Table LABEL:minedArchitecturalInfo). For example, Liu et al. [S63] utilized a semi-automatic approach for mining architectural pattern use scenarios and related pattern pairs from specific software development blogs and Q&A site (i.e., Stack Overflow). This mined architectural information can aid developers to acquaint with the usage and relatedness of architecture patterns in today’s modern software development. On the other hand, Chinnappan et al. [S9] employed a manual approach (i.e., qualitative data analysis) to mine energy-aware architectural tactics (such as energy savings mode tactic, stop current task & recharge tactic) for Robotic Operating System (ROS) based software systems from GitHub, ROS Answers, and Stack Overflow. To do so, they carried out a multi-phase study that resulted in seven energy-awareness architectural tactics. To foster the applicability of the identified tactics even beyond the ROS community, they described these tactics in a generic implementation-independent manner by utilizing diagrams inspired by UML component and sequence diagram notations. These energy-aware architectural tactics can serve as guidance for roboticists, as well as other developers interested in architecting and implementing energy-aware software. Then again, Márquez et al. [S26] proposed a semi-automatic approach for mining frameworks in OSS for addressing scalability concerns in microservice applications, and these frameworks can assist developers in designing scalable microservice applications.
@@
Architectural decisions: Mining and understanding architectural decisions can help inform future architectural decisions and implementation choices, and can avoid introducing architectural inefficiencies (Shahbazian et al., 2018) or architectural decay later (Shahbazian et al., 2018). Thus, 32 out of 104 studies (30.8%) focused on mining architectural decisions to assist the development. We categorized the mined architectural decisions into four subcategories (including structural decision, behavioral decision, and technology decision) by following existing classifications of architectural design decisions in (Kruchten, 2004) and (Bass et al., 2012) (see Table LABEL:minedArchitecturalInfo). For example, Bhat et al. [S22] employed an ML-based approach to mine and classify architectural decisions from issue tracking systems (e.g., Jira) into three subcategories, namely structural, behavioral, and non-existence decisions. The mined architectural decisions in this category can help architects and developers utilize these decisions in similar development contexts. Figure 6(a) illustrates an example of mined architectural decisions (i.e., structure decisions) from an issue tracking system.
@@
System requirements: Analyzing requirements to relate and map them to their corresponding architectural elements (e.g., architectural components) has become one of the major challenges faced by architects during the development process (Casamayor et al., 2012a). The failure of a high percentage of software projects is often caused by, for example, the lack of proper requirements analysis, incomplete requirement specifications, and changing requirements, among others (Hull et al., 2005). 31 out of 104 studies (29.8%) focus on mining the requirements of systems from repositories to assist architecting process (see Table LABEL:minedArchitecturalInfo). We further categorized the mined system requirements into three subcategories: Quality Attributes (QAs), functional requirement, and constraint by following the classification described in (Cervantes and Kazman, 2016). For instance, in [S66], Gokyer et al. used an approach based on NLP and ML techniques for automatically mining Non-Functional Requirements (NFRs) expressed in plain text and mapping these NFRs to architectural elements, such as architectural components, in the problem domain. The mined architectural information can guide architects in making architectural decisions effectively. On the other hand, in [S43], Casamayor et al. presented an approach based on NLP and ML techniques to semi-automatically mine and analyze functional requirements (from the textual description of requirements) that will become the responsibilities of certain architectural components in the system, in order to help bridge the gap between requirements analysis and architectural design. Then again Soliman et al. [S3] developed a Web-based search engine for searching several types of architectural information (including constraints) from Q&A sites.